Evaluate if possible ^3/-64 [cube root of -64] and ^3/64 [cube root of 64] and /-64 [square root (-64)] and /64 [square root (64)]
Answer by RAY100(1637)

cube root of (-64) = -4,,,,,(-4)^3 = -64
.
cube root of (64) = 4,,,,,,,(4)^3 =64
.
sqrt (-64) = sqrt(64)*sqrt(-1) = 8i,,,,,,,(8i)^2 = 64i^2 =-64
.
remember sqrt(-1) =i,,,,,,i^2 =-1
.
sqrt(64) =8,,,,,,,,,,,,,,,,,,,,,,,,,,,,8^2 =64
.